MY PAPER TITLE

?First author name

?Organization name

?Organization address ?Email: someone@some.place

?Second author name

?Organization name

?Organization address ?Email: someoneelse@some.place

Keywords: Resource allocation, large-scale, decoupling, optimization.

Abstract:  Many Internet service providers and online services require you to manually enter information, such as your user name and password, to establish a connection. With Scripting support for Dial-Up Networking, you can write a script to automate this process. A script is a text file that contains a series of commands, parameters, and expressions required by your Internet service provider or online service to establish the connection and use the service. You can use any text editor, such as Microsoft Notepad, to create a script file. Once you’ve created your script file, you can then assign it to a specific Dial-Up Networking connection by running the Dial-Up Scripting Tool. Many Internet service providers and online services require you to manually enter information, such as your user name and password, to establish a connection. With Scripting support for Dial- Up Networking, you can write a script to automate this process. A script is a text file that contains a series of commands, parameters, and expressions required by your Internet service provider or online service to establish the connection and use the service. You can use any text editor, such as Microsoft Notepad, to create a script file. Once you’ve created your script file, you can then assign it to a specific Dial-Up Networking connection by running the Dial-Up Scripting Tool.

1 INTRODUCTION

Many Internet service providers and online services require you to manually enter information, such as your user name and password, to establish a connection. With Scripting support for Dial-Up Networking, you can write a script to automate this process. Many Internet service providers and online services (Smith, 1998) require you to manually enter information, such as your user name and password, to establish a connection. With Scripting support for Dial-Up Networking, you can write a script to automate this process. Many Internet service providers and online services (Louis, 1999) require you to manually enter information, such as your user name and password, to establish a connection. With Scripting support for Dial-Up Networking, you can write a script to automate this process.

2 BASIC STRUCTURE OF A SCRIPT

A command is the basic instruction that a script file contains. Some commands require parameters that further define what the command should do. An expression is a combination of operators and arguments that create a result. Expressions can be used as values in any command. Examples of expressions include arithmetic, relational comparisons, and string concatenations. A command is the basic instruction that a script file contains. Some commands require parameters that further define what the command should do. An expression is a combination of operators and arguments that create a result. Expressions can be used as values in any command. Examples of expressions include arithmetic, relational comparisons, and string concatenations.

A command is the basic instruction that a script file contains. Some commands require parameters that further define what the command should do. An expression is a combination of operators and arguments that create a result. Expressions can be used as values in any command. Examples of expressions include arithmetic, relational comparisons, and string concatenations.

The basic form of a script for Dial-Up Networking follows:

  ; A comment begins with a semi-colon
  ;and extends to the end of the line.
  proc main
  ; A script can have any number of
  ; variables and commands
  variable declarations
  command block
  endproc

You must declare variables before you add commands. The first command in the main procedure is executed, and then any subsequent commands are executed in the order they appear in the script. The script ends when the end of the main procedure is reached.

Figure 1: Example of a figure.

3 VARIABLES

Scripts may contain variables. Variable names must begin with a letter or an underscore, and may contain any sequence of upper- or lower-case letters, digits, and underscores. You cannot use a reserved word as a variable name. For more information, see the list of reserved words at the end of this document. Variables are assigned values using the following assignment statement:

variable = expression  (1)

The variable gets the evaluated expression.

Examples:

  integer count = 5

                                                    integer timeout = (4 * 3)
                                                    integer i
                                                    boolean bDone = FALSE
                                                    string szIP = (getip 2)
                                                    set ipaddr szIP

3.1 System Variables

System variables are set by scripting commands or are determined by the information your enter when you set up a Dial-Up Networking connection.

System variables are read-only, which means they cannot be changed within the script.

4 CONCLUSION

All the conclusions and future work in here.

REFERENCES

Louis, R. (1999). Software agents activities. In ICEIS’99, 1st International Conference on Enterprise Information Systems. ICEIS Press.

Smith, J. (1998). The Book. The publishing company, London, 2nd edition.